OpenStudy (anonymous):

Which shows solutions to x + 4y = 9 if x and y must be whole numbers. A. {(1, 2), (5, 1), (9,0)} B. {(0, 9), (1, 4), (2, 1)} C. {(2, 1), (1, 6), (0, 9)} D. {(0, 9), (1, 5), (2, 1)}

OpenStudy (anonymous):

hey mikayla

OpenStudy (anonymous):

HEYY xD

OpenStudy (anonymous):

brad!!!!:D

OpenStudy (anonymous):

hows it goin

OpenStudy (anonymous):

A 1+4(2)=9 5+4(1)=9 9+4(0)=9

OpenStudy (anonymous):

1+8=9 5+4=9 9+0=9